A stationary billiard ball, with a mass of 0.17 kg, is struck by an identical ball moving at 4.0 m/s. After the collision, the second ball moves 60.09 to the left of its original direction. The stationary ball moves 30.0° direction. What is the velocity of each ball after the collision? 24. to the right of the moving balls original 25. A 1345-kg car moving east at 15.7 m/s is struck by a 1923-kg car moving north. They 63.5o. Was the are stuck together and move with an initial velocity of 14.5 m/s at north-moving car exceeding the 20.1 m/s speed limit?

A stationary billiard ball with a mass of 0.17kg, is sturck by an identical ball moving at 4.0 m/s. After the collision, the second ball moves 60 degrees to the left of its original direction. The stationary ball movies 30 degrees to the right of the moving ball's original direction. What is the velocity of each ball after the collision?
Answer #1

Here ,

let the velocities of the balls after the collision is v1 and v2

Now, Using conservation of momentum along the line of collision

0.17 * 4 = 0.17 * v1 * cos(60) + 0.17 * v2 * cos(30)

using conservation of momentum perpendicular to line of collision

0.17 * v1 * sin(60) - 0.17 * v2 * sin(30) = 0

solving for v1 and v2

v1 = 2 m/s

v2 = 3.46 m/s

the velocities of the balls are 2 m/s and 3.46 m/s